OpenStudy (anonymous):

how do i factor: x^5 + y^5

OpenStudy (anonymous):

similar to the sum of two cubes start with \((x+y)\times \text{something}\)

OpenStudy (anonymous):

remember that an exponent times an exponent is the sum of the exponents

OpenStudy (dumbcow):

https://www.wolframalpha.com/input/?i=Factor+x^5+%2By^5 sorry im lazy

OpenStudy (anonymous):

probably easiest to remember that \[a^5-b^5=(a-b)(a^4+a^3b+a^2b^2+ab^3+b^4)\]nice easy pattern, then replace \(a\) by \(x\) and \(b\) by \(-y\) be careful of your signs

OpenStudy (anonymous):

@dumbcow if you were really lazy, you wouldn't have written "factor" because the wolf will do it anyway

OpenStudy (dumbcow):

haha true but its not as clear..you you have to go down and find "alternate form"